OpenStudy (bloomlocke367):

A jar contains 7 blue cubes, 4 blue spheres, 5 green cubes, and 6 green spheres. If you select an object at random, what is the object is green or a cube?

OpenStudy (bloomlocke367):

is it 23/22?

OpenStudy (bloomlocke367):

@FateRocks8D

OpenStudy (anonymous):

Add all the stuff up: 7 + 4 + 5 + 6 = 11 + 11 = 22 probability of a green object or a cube, so add the green objects and the total cubes 11 green objects + 12 cubes. = 23 23 / 22

OpenStudy (bloomlocke367):

Okay, that's what I did.. So I'm right?

OpenStudy (bloomlocke367):

Nope, I was wrong. It's 9/11... how is it 9/11?

OpenStudy (bloomlocke367):

@FateRocks8D

OpenStudy (anonymous):

Lol wasn't finised my computor froze

OpenStudy (bloomlocke367):

Oh. Dang. Well what's next?

OpenStudy (anonymous):

so 22 total objects - 4 blue spheres = 18 / 22 = 9 / 11

OpenStudy (bloomlocke367):

so, why'd we add the cubes and the green objects?

OpenStudy (anonymous):

mhmm and now ur answer is 9/11 :) wow u did better then me good job

OpenStudy (bloomlocke367):

I'm lost, still

OpenStudy (anonymous):

if u go back to the stuff I just touch u I don't think u wouldn't :P Pay attaion tag me in ur questions imam help someone else

OpenStudy (bloomlocke367):

What I don't understand is why we added the green objects and the cubes if we could just subtract the blue spheres... @FateRocks8D

OpenStudy (anonymous):

idk I thought it wold be faster

OpenStudy (bloomlocke367):

oh, okay